OpenStudy (anonymous):

a softball pitcher throws a pitch horizontally towards home plate at 24 m/s. if home plate is 13.1 m away calculate how far the ball will drop as it travels towards home.

OpenStudy (anonymous):

The drop of the ball will only be due to vertical forces. In this case, the only force acting will be gravity. If we know how long the ball will be in the air, we can figure out how far it will drop in that time. To find the time taken to reach the catcher, we can divide the distance by the velocity. t = x / v = 13.1 / 24 = 0.546 s So, we can calculate the amount the ball will fall in this time using the free fall equation: y = (1/2) g t^2 y = (1/2) (9.8) (0.546)^2 y = 1.46 m Therefore, the ball will fall 1.46 m on it's way to home plate.